# Low-energy interactions between doubly charmed baryons and Goldstone bosons from lattice QCD

### v1

@@ 4.3 Pole analysis of the scattering amplitude
The interaction in the $\Xi_{cc}K^{(1,0)}$ channel is attractive, manifesting as negative energy shifts in figure 1. A near-threshold pole at $p^{2}\approx-0.03\ \mathrm{GeV}^{2}$ is discovered in the upper half-plane for both $M_{\pi}\sim 210$ and $300$ MeV, as shown in figure 2(c). This pole suggests the existence of a possible virtual state and indicates that the strength of the attraction is too weak to form a bound state. However, the stability of this pole against ERE truncation remains unevaluated due to insufficient data for performing NLO fits.
@@ Acknowledgements.
We would like to thank Ying Chen, Feng-Kun Guo, Qu-Zhi Li, Chuan Liu, Peng Sun, Xiao-Nu Xiong and Yi-Bo Yang for helpful discussions. This work is supported by Science Research Project of Hebei Education Department under Contract No. QN2025063; National Nature Science Foundations of China (NSFC) under Contract Nos. 12275076, 11905258, 12335002, 12175279, 12293060, 12293061; by the Science Fund for Distinguished Young Scholars of Hunan Province under Grant No. 2024JJ2007; by Hebei Natural Science Foundation under Grant No. A2025205018; by the Science Foundation of Hebei Normal University with Contract No. L2025B09.
@@ Appendix A Extraction of energies
Appendix B Influence of coupled-channel effects
As mentioned in section 3.3, we must consider some channels that may be encountered in our analysis of the two-particle system. The operators of $D^{+}$, $D^{0}$, $D_{s}^{+}$, $\Lambda_{c}^{+}$, $\Sigma_{c}^{++}$, and $\Omega_{c}^{0}$ are needed, which are constructed as
@@ Appendix B Influence of coupled-channel effects
We perform simulations of the two-point correlation functions with the above interpolating fields. The corresponding masses are obtained and are listed in table 7. The influence of the $B_{c}D$ channels on our energy levels of $B_{cc}\phi$ systems is illustrated in figure 1, where $B_{c}\in\{\Sigma_{c}^{++},\Omega_{c}^{0},\Lambda_{c}^{+}\}$ and $D\in\{D^{+},D^{0},D_{s}^{+}\}$.

### v2

@@ 4.3 Pole analysis of the scattering amplitude
The interaction in the $\Xi_{cc}K^{(1,0)}$ channel is attractive, manifesting as negative energy shifts in figure 1. From the perspective of HDA symmetry, this channel corresponds to the $I=0$ channel of $D\bar{K}$ scattering. Previous studies have reported a virtual state in the corresponding $D\bar{K}$ channel Guo:2015dha ; Cheung:2020mql . Motivated by this, we investigate the pole position in the $\Xi_{cc}K^{(1,0)}$ system to explore the potential existence of a counterpart state. A near-threshold pole at $p^{2}\approx-0.03\ \mathrm{GeV}^{2}$ is discovered in the upper half-plane for both $M_{\pi}\sim 210$ and $300$ MeV, as shown in figure 2(c). This pole suggests the existence of a possible virtual state and indicates that the strength of the attraction is too weak to form a bound state. However, the stability of this pole against ERE truncation remains unevaluated due to insufficient data for performing NLO fits.
@@ Acknowledgements.
We would like to thank Ying Chen, Feng-Kun Guo, Qu-Zhi Li, Chuan Liu, Peng Sun, Xiao-Nu Xiong and Yi-Bo Yang for helpful discussions. We thank the CLQCD collaboration for providing us their gauge configurations with dynamical fermions CLQCD:2023sdb ; CLQCD:2024yyn , which are generated on the Southern Nuclear Science Computing Center (SNSC), HPC Cluster of ITP-CAS, and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. This work is supported by Science Research Project of Hebei Education Department under Contract No. QN2025063; National Natural Science Foundations of China (NSFC) under Contract Nos. 12275076, 11905258, 12335002, 12175279, 12293060, 12293061; by the Science Fund for Distinguished Young Scholars of Hunan Province under Grant No. 2024JJ2007; by Hebei Natural Science Foundation under Grant No. A2025205018; by the Science Foundation of Hebei Normal University with Contract No. L2025B09.
@@ Appendix A Extraction of energies
Appendix B Details of the interpolating operators
where $C=i\gamma_{2}\gamma_{0}$ is the charge conjugation matrix and the positive parity projector is defined by $P_{+}=(1+\gamma_{0})/2$. For the two-particle operators, the coefficients $C_{\alpha,\mathbf{p}_{\mathbf{1}},\mathbf{p}_{\mathbf{2}}}$ in eq. (3) are listed in Table 6.
$p=0$ 11 $(0,0,0)$ $(0,0,0)$ 11
$p=1$ 1 (1,0,0)(-1,0,0) $(1,0,0)$ 11
11 $(1,0,0)$ (1,0,0)(-1,0,0) 11
11 (0,1,0)(0,-1,0) $(0,1,0)$ 11
11 $(0,1,0)$ (0,1,0)(0,-1,0) 11
11 (0,0,1)(0,0,-1) $(0,0,1)$ 11
11 $(0,0,1)$ (0,0,1)(0,0,-1) 11
$p=\sqrt{2}$ 11 (1,1,0)(-1,-1,0) $(1,1,0)$ 11
11 $(1,1,0)$ (1,1,0)(-1,-1,0) 11
11 (1,0,1)(-1,0,-1) $(1,0,1)$ 11
11 $(1,0,1)$ (1,0,1)(-1,0,-1) 11
11 (0,1,1)(0,-1,-1) $(0,1,1)$ 11
11 $(0,1,1)$ (0,1,1)(0,-1,-1) 11
11 (1,1,0)(-1,1,0) (1,1,0)(1,-1,0) 11
11 (1,1,0)(1,-1,0) (1,1,0)(-1,1,0) 11
11 (1,0,1)(-1,0,1) (1,0,1)(1,0,-1) 11
11 (1,0,1)(1,0,-1) (1,0,1)(-1,0,1) 11
11 (0,1,1)(0,1,-1) (0,1,1)(0,-1,1) 11
11 (0,1,1)(0,-1,1) (0,1,1)(0,1,-1) 11
Furthermore, as mentioned in section 3.3, we must consider some channels that may be encountered in our analysis of the two-particle system. The operators of $D^{+}$, $D^{0}$, $D_{s}^{+}$, $\Lambda_{c}^{+}$, $\Sigma_{c}^{++}$, and $\Omega_{c}^{0}$ are needed, which are constructed as
@@ 11 $(0,1,1)$ (0,1,1)(0,-1,-1) 11
We perform simulations of the two-point correlation functions with the above interpolating fields. The corresponding masses are obtained and are listed in table 7. The thresholds of $B_{c}D$ channels are illustrated in figure 1.